In an experiment to determine the thermal conductivity of a bar of new alloy, one end of the bar is maintained at 0 °C and the other end at 100 °C. The bar has a cross-sectional area of 1.0 cm2 and a length of 15 cm. If the rate of heat conduction through the bar is



34 W, what is the thermal conductivity of the bar?





1
Expert's answer
2021-12-07T18:48:20-0500

"\frac{Q}{\Delta\>t}=\frac{K A\Delta\>T}{\Delta\>x}"


"34=\frac{K(1×10^{-4})×(100-0)}{15×10^{-2}}"


"K=510w\>m^{-1}K^{-1}"